if 65g of ice at 0.0 degrees celsius is placed in 245mL of water at 35 degrees celsius, find the final temperature if all of the ice melts.

STEP BY STEP PLEASE
Answer #1

Latent heat of melting of ice = 334 J/g (The amount of heat required to melt 1 gm of ice)

Thus, total heat require to melt 65 gms of ice = 334*65 = 21710 joules

Specific heat of water = 4.186 joules/gram C

In given condition, heat given by the water to melt the ice = 245 * 4.186 * (Ti - Tf)

Assuming density of water = 1g/cc

Equating the equations :

21710 = 245 * 4.186 * (Ti-Tf)

Final Temperature = 13.83 C